OpenStudy (anonymous):

The equation A = x(x - 2) describes the area, A, of a rectangular flower garden, where x is the width in feet. Medal What would be the width in feet of the flower garden if the area is 15 square feet?

OpenStudy (danjs):

set up an equation

OpenStudy (danjs):

Area = Length x Width

OpenStudy (danjs):

x = width so x-2 = Length

OpenStudy (danjs):

Total area = 15 15 = x(x-2)

OpenStudy (danjs):

I would put that into a form of Ax^2 + Bx + C = 0 and try to factor it

OpenStudy (danjs):

x^2 - 2x - 15 = 0

OpenStudy (danjs):

(x+3)(x-5) = 0 so (x - 5) = 0 or (x+3) = 0

OpenStudy (anonymous):

So what will be the answer out of these answers.. A.)3 B.)6 C:)5 D.)2

OpenStudy (danjs):

So if the Width is X, what is the answer

OpenStudy (danjs):

What are the solutions to x - 5 = 0 or x + 3 = 0

OpenStudy (danjs):

x is a measure of width, a length, so it should be the positive of the two values

OpenStudy (anonymous):

Would it be 5?

OpenStudy (danjs):

yes

OpenStudy (danjs):

Try that in your original given equation of Area = x(x-2) For x = 5 and see if you get 15, as a check

OpenStudy (danjs):

5(5-2) = 5*3 = 15 good